# QOCO: A Quadratic Objective Conic Optimizer with Custom Solver Generation

### v2

@@ Appendix B Custom LDL Factorization Performance
$\begin{split}\underset{x,u}{\mathrm{minimize}}\quad&\frac{1}{2}\left(\sum_{k=1}^{T}x_{k}^{\top}Q_{k}x_{k}+\sum_{k=1}^{T-1}u_{k}^{\top}R_{k}u_{k}\right)\\ \mathrm{subject\;to}\quad&x_{k+1}=A_{k}x_{k}+B_{k}u_{k}\;\quad\forall k\in[1,T-1]\\ &x_{1}=x_{\mathrm{init}},\\ \end{split}$
@@
$\begin{split}P&=\mathrm{blkdiag}(Q_{1},Q_{2},\ldots,Q_{T},R_{1},R_{2},\ldots,R_{T-1})\\ H&=\begin{bmatrix}A_{1}&-I&&&&&B_{1}\\ &A_{2}&-I&&&&&B_{2}\\ &&&\ddots&&&&&\ddots\\ &&&&A_{N-1}&-I&&&&B_{N-1}\\ I&&&&&&&&&\\ \end{bmatrix}.\end{split}$
@@ C.1 Robust Kalman Filter
$\begin{split}\underset{x_{k},w_{k},v_{k}}{\text{minimize}}\quad&\sum_{k=0}^{N-1}\left(\|w_{k}\|_{2}^{2}+\tau\phi_{\rho}(v_{k})\right)\\ \text{subject to}\quad&x_{k+1}=Ax_{k}+Bw_{k}\quad\forall k\in[0,N-1]\\ \quad&y_{k}=Cx_{k}+v_{k}\quad\forall k\in[0,N-1],\\ \end{split}$ (33)
@@
$\phi_{\rho}(z)=\begin{cases}\|z\|_{2}^{2}&\|z\|_{2}\leq\rho\\ 2\rho\|z\|_{2}-\rho^{2}&\|z\|_{2}>\rho.\\ \end{cases}$
In this problem, $x_{k}\in\mathbb{R}^{4}$ represents the state of the vehicle at timestep $k$, $w_{k}\in\mathbb{R}^{2}$ is an unknown force vector acting on the vehicle at timestep $k$, $v_{k}\in\mathbb{R}^{2}$ is the measurement noise vector at timestep $k$, and $y_{k}\in\mathbb{R}^{2}$ represents a noisy measurement of the vehicle’s state at timestep $k$. The matrices $A$ and $B$ are known dynamics matrices for the system, and $C$ is the known observation matrix. To reconstruct the vehicle’s state trajectory, we solve Problem 33.
@@
$\vspace{10mm}B$ $=\begin{bmatrix}\frac{1}{2}\Delta t^{2}&0&\\ 0&\frac{1}{2}\Delta t^{2}\\ \Delta t&0\\ 0&\Delta t\end{bmatrix}$
@@
[…]
@@ C.2 Lossless Convexification
$\vspace{10mm}B$ $=\begin{bmatrix}\frac{1}{2}\Delta t^{2}(I_{3})\\ \Delta t(I_{3})\end{bmatrix}$
@@
$g_{0}=9.807,\;\rho_{1}=100,\;\rho_{2}=500,\;m_{\mathrm{dry}}=25,\;m_{\mathrm{wet}}=35,\;\theta_{\mathrm{max}}=\pi/4,\;\alpha=0.001.\$
@@ C.4 Portfolio Optimization
$\begin{split}\underset{x,y}{\text{minimize}}\quad&x^{\top}Dx+y^{\top}y-\gamma^{-1}\mu^{\top}x\\ \text{subject to}\quad&y=F^{\top}x\\ \quad&\sum_{i=1}^{n}x_{i}=1\\ \quad&x_{i}\geq 0\quad\forall i\in[1,n].\\ \end{split}$ (34)
@@ C.5 Oscillating Masses
$\vspace{10mm}B_{c}$ $=\begin{bmatrix}0_{N\times N}\\ I_{N}\end{bmatrix}.$
Here, $L_{N}\in\mathbb{R}^{N\times N}$ is a symmetric tridiagonal matrix with $-2$ on the main diagonal and $1$ on the subdiagonal and superdiagonal.
@@
For each problem size (i.e., each number of timesteps), we generate 20 random problem instances. This involves generating a random initial state $x_{\mathrm{init}}$, as well as the matrices $Q$ and $R$. We draw each element of $x_{\mathrm{init}}$ from $\mathcal{N}(0,1)$, then clip the elements such that they lie within the interval $[-0.9x_{\mathrm{max}},0.9x_{\mathrm{max}}]$. This ensures that the initial state satisfies the state constraints, making the resulting problem feasible. The matrices $Q$ and $R$ are generated as diagonal matrices, with each diagonal element drawn from the uniform distribution $\mathcal{U}(0,10)$.

### v3

@@ Appendix B Custom LDL factorization performance
[…]
@@
[…]
@@ C.1 Robust Kalman filter
[…]
@@
[…]
In this problem, $x_{k}\in\mathbb{R}^{4}$ represents the state of the vehicle at timestep $k$, $w_{k}\in\mathbb{R}^{2}$ is an unknown force vector acting on the vehicle at timestep $k$, $v_{k}\in\mathbb{R}^{2}$ is the measurement noise vector at timestep $k$, and $y_{k}\in\mathbb{R}^{2}$ represents a noisy measurement of the vehicle’s state at timestep $k$. The matrices $A$ and $B$ are known dynamics matrices for the system, and $C$ is the known observation matrix. To reconstruct the vehicle’s state trajectory, we solve Problem 32.
@@
B
[…]
$=\begin{bmatrix}\frac{1}{2}\Delta t^{2}&0&\\ 0&\frac{1}{2}\Delta t^{2}\\ \Delta t&0\\ 0&\Delta t\end{bmatrix}$
@@
For each problem size, we generate 20 random problem instances, corresponding to 20 different observation trajectories $y_{t}$. To do this, we first randomly generate the force and noise vectors $w_{t}$ and $v_{t}$, respectively, for $t\in[0,N-1]$. Then, we forward-simulate an initial state of $x_{0}=[0\;0\;0\;0]$ through the dynamics defined in the constraints of Problem 32. Each element of $w_{t}$ and $v_{t}$ is drawn from a standard Gaussian distribution $\mathcal{N}(0,1)$. Additionally, for 20 % of the elements of $v_{t}$, we introduce outlier noise by drawing elements from $\mathcal{N}(0,20)$, to simulate noise with a larger magnitude.
@@ C.2 Lossless convexification
B
@@
$=\begin{bmatrix}\frac{1}{2}\Delta t^{2}(I_{3})\\ \Delta t(I_{3})\end{bmatrix}$
$g$ $=[0\;\;0\;\;-0.5g_{0}\Delta t^{2}\;\;0\;\;0\;\;-g_{0}\Delta t].$
@@
$g_{0}=9.807,\;\rho_{1}=100,\;\rho_{2}=500,\;m_{\mathrm{dry}}=25,\;m_{\mathrm{wet}}=35,\;\theta_{\mathrm{max}}=\pi/4,\;\alpha=0.001.\\$
@@ C.4 Portfolio optimization
[…]
@@ C.5 Oscillating masses
Bc
[…]
$=\begin{bmatrix}0_{N\times N}\\ I_{N}\end{bmatrix}$.
Here, $L_{N}\in\mathbb{R}^{N\times N}$ is a symmetric tridiagonal matrix with 2-2 on the main diagonal and $1$ on the subdiagonal and superdiagonal.
@@
For each problem size (i.e., each number of timesteps), we generate 20 random problem instances. This involves generating a random initial state $x_{\mathrm{init}}$, as well as the matrices $Q$ and $R$. We draw each element of $x_{\mathrm{init}}$ from $\mathcal{N}(0,1)$, then clip the elements such that they lie within the interval $0.9x_{\mathrm{max}},0.9x_{\mathrm{max}}]$. This ensures that the initial state satisfies the state constraints, making the resulting problem feasible. The matrices $Q$ and $R$ are generated as diagonal matrices, with each diagonal element drawn from the uniform distribution $\mathcal{U}(0,10)$.